# On Selection of Cross-Section Averages in Non-stationary Environments

### v2

@@ 6.4 Part $M_{0}\subset M$
We now move to the case when $M_{0}\subset M$ (over-specification). To analyze this case we will introduce additional notation. We decompose the selector matrix $\mathbf{q}_{M}=[\mathbf{q}_{M_{0}},\mathbf{q}_{M_{0}^{C}}]$, where $\mathbf{q}_{M_{0}}\in\mathbb{R}^{k\times m}$ and $\mathbf{q}_{M_{0}^{C}}\in\mathbb{R}^{k\times(g-m)}$, where the latter selector corresponds to $g-m$ excess (over-selected) averages. By using this, we can further decompose
@@
where $\left\|\overline{\mathbf{V}}_{M_{1}}^{0}\right\|=O_{p}(N^{-1/2})$, but $\left\|\overline{\mathbf{V}}_{M_{2}}^{0}\right\|=O_{p}(1)$, such that the last $g-m$ columns are non-degenerate.
@@
where $\mathbf{S}=\mathrm{diag}(T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}},T^{-1}\overline{\mathbf{V}}_{M_{2}}^{0\prime}\overline{\mathbf{V}}_{M_{2}}^{0})\in\mathbb{R}^{g\times g}$. The rate comes from the leading, but vanishing components of the last three terms on the right-hand side of (6.4). Specifically,
@@
by Corollary (1). This implies that
$\left\|(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}-\mathbf{S}^{+}\right\|=O_{p}(N^{-1/2}T^{-\tau/2})+O_{p}(T^{-1/2}),$ (A.49)
@@
Hence, if we insert (6.4) into $\mathbf{II}$, we obtain $\mathbf{II}=\mathbf{II}^{1}-\mathbf{II}^{2}-\mathbf{II}^{3}+\mathbf{II}^{4}$, and so
@@
$+\underbrace{\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1/2}\mathbf{V}_{i}^{\prime}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}\right\|^{2}\left\|(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}-\mathbf{S}^{+}\right\|}_{=o_{p}(C_{N,T})^{-2}}$
@@
and so the last term in (6.4) is $o_{p}(C_{N,T}^{-2})$. The other terms are clearly dominated. Next, we go to
@@
and so the overall order is dominated by the first and third terms in (6.4). Further, note that $\mathbf{II}^{3}$ is just a transpose of $\mathbf{II}^{2}$ and so
@@
$\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})$ $=\underbrace{\ln\mathrm{det}\left(\mathbf{I}_{k}+\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right]\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)}_{\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}}+k(g-m)p_{N,T}$
@@
for some integer $n$. However, given that $p_{N,T}O_{p}(C_{N,T}^{-2})\to\infty$, then we have that
@@
$\mathbb{P}\left(\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})<0\right)=\mathbb{P}\left(p_{N,T}^{-1}(\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0}))<0)\right)\to 0,$ (A.77)
@@
6.5 Proof Changes for IC of Margaritella and Westerlund, (2023)
Note that IC of MW is based on the residuals from the first stage CCE regression under the full set of available $k+1$ CAs. In particular, let us assume homogeneous $\boldsymbol{\beta}$. Then
@@ 6.5 Proof Changes for IC of Margaritella and Westerlund, (2023)
$\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})$ $=\underbrace{\ln\left[1+\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}\right)\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}\widehat{\boldsymbol{\nu}}_{i}\right)^{-1}\right]}_{\mathcal{V}_{MW}-\mathcal{V}_{MW}^{0}}$
$+(g-m)\cdot p_{N,T},$ (A.79)
@@
$\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}$ $=\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})$
$-\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
$-\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})$
$+\frac{1}{NT}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}).$ (A.80)
@@
$\left\|\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$ $\leq(NT)^{-1}\underbrace{\left\|\sqrt{NT}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|^{2}}_{O_{p}(1)}\underbrace{\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|}_{O_{p}(T^{\tau})}$
=$=\begin{cases}O_{p}(N^{-1}T^{\tau-1})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.81)
because the rate of $\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|$ was already derived for the IC of DVS under both cases. Also, $\left\|\sqrt{NT}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|=O_{p}(1)$ under general factors (see Westerlund,, 2018), and thus the $\sqrt{NT}$-consistency will hold under our assumptions, as well. Similarly,
@@
$\left\|\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})\right\|$ $\leq(NT)^{-1/2}\left\|\sqrt{NT}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$
@@
×$\times\underbrace{\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})\right\|}_{O_{p}(T^{\tau})}$
=$=\begin{cases}O_{p}(N^{-1/2}T^{\tau-1/2}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.82)
@@
because asymptotically $\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$ behaves identically to $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$, and thus the dominating component will have the same rate. Eventually,
[…]
which will be the dominating term.

### v3

@@ 5.3.2 Case of $M_{0}\subset M$
We now move to the case when $M_{0}\subset M$ (over-specification). To analyze this case we will introduce additional notation. We decompose the selector matrix $\mathbf{q}_{M}=[\mathbf{q}_{M_{0}},\mathbf{q}_{M_{0}^{C}}]$, where $\mathbf{q}_{M_{0}}\in\mathbb{R}^{k\times m}$ and $\mathbf{q}_{M_{0}^{C}}\in\mathbb{R}^{k\times(g-m)}$, where the latter selector corresponds to gmg-m excess (over-selected) averages. By using this, we can further decompose
@@
where $\left\|\overline{\mathbf{V}}_{M_{1}}^{0}\right\|=O_{p}(N^{-1/2})$, but $\left\|\overline{\mathbf{V}}_{M_{2}}^{0}\right\|=O_{p}(1)$, such that the last gmg-m columns are non-degenerate.
@@
where $\mathbf{S}=\mathrm{diag}(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}},T^{-1}\overline{\mathbf{V}}_{M_{2}}^{0\prime}\overline{\mathbf{V}}_{M_{2}}^{0})\in\mathbb{R}^{g\times g}$. The rate comes from the leading, but vanishing components of the last three terms on the right-hand side of (5.3.2). Specifically,
@@
by Corollary (1). By following the approach in Bai, (2004), we see that this implies that
$\left\|(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}-\mathbf{S}^{+}\right\|$ $=\left\|\mathbf{S}^{+}(\mathbf{S}-\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}\right\|$
$\leq\left\|\mathbf{S}^{+}\right\|\left\|\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}-\mathbf{S}\right\|\left\|(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}\right\|$
$=O_{p}(N^{-1/2}T^{-\tau/2})+O_{p}(T^{-1/2}),$ (A.53)
@@
Hence, if we insert (5.3.2) into $\mathbf{II}$, we obtain $\mathbf{II}=\mathbf{II}^{1}-\mathbf{II}^{2}-\mathbf{II}^{3}+\mathbf{II}^{4}$, and so
@@
$+\underbrace{\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1/2}\mathbf{V}_{i}^{\prime}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1}\right\|^{2}\left\|(\mathbf{G}_{T}^{-1}\widehat{\mathbf{F}}^{0\prime}_{M_{\mathbf{x}}}\widehat{\mathbf{F}}^{0}_{M_{\mathbf{x}}}\mathbf{G}_{T}^{-1})^{+}-\mathbf{S}^{+}\right\|}_{=o_{p}(C_{N,T}^{-2})}$
@@
and so the last term in (5.3.2) is $o_{p}(C_{N,T}^{-2})$. The other terms are clearly dominated. Next, we go to
@@
and so the overall order is dominated by the first and third terms in (5.3.2). Further, note that $\mathbf{II}^{3}$ is just a transpose of $\mathbf{II}^{2}$ and so
@@
for some integer $n$. However, given that $p_{N,T}^{-1}O_{p}(C_{N,T}^{-2})=o(1)$, then we have that
@@
$\mathbb{P}\left(\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0})<0\right)=\mathbb{P}\left(p_{N,T}^{-1}(\mathrm{IC}^{DVS}(M)-\mathrm{IC}^{DVS}(M_{0}))<0\right)\to 0,$ (A.81)
@@
5.3.3 Proof Changes for IC of Margaritella and Westerlund, (2023)
5.3.3.1  Homogeneous $\boldsymbol{\beta}$
Note that IC of MW is based on the residuals from the first stage CCE regression under the full set of available $k+1$ CAs (note that $|M|=k+1$ now). In this case, we have $\widehat{\mathbf{F}}=[\overline{\mathbf{y}},\overline{\mathbf{X}}]=\mathbf{F}\overline{\mathbf{C}}+\overline{\mathbf{U}}$, where now $\overline{\mathbf{C}}$ is a loading matrix with rank of $m$ even when $N\to\infty$. In particular, let us assume homogeneous $\boldsymbol{\beta}$. Then
@@ 5.3.3.1  Homogeneous $\boldsymbol{\beta}$
$\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})$ $=\ln\left[1+\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}\right)\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}\widehat{\boldsymbol{\nu}}_{i}\right)^{-1}\right]$
$+(g-m)\cdot p_{N,T}$
$=\ln\left[1+T^{\tau}\left(\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}\right)\left(\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}\widehat{\boldsymbol{\nu}}_{i}\right)^{-1}\right]$
$+(g-m)\cdot p_{N,T}$ (A.83)
@@
$\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}$ $=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})$
$-\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
$-\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})$
$+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}).$ (A.84)
@@
$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$ (A.85)
@@
$\leq(NT)^{-1}\underbrace{\left\|\sqrt{NT}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|^{2}}_{O_{p}(1)}\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|$
$=\begin{cases}O_{p}((NT)^{-1})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.86)
because the rate of $\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|$ was already derived for the IC of DVS under both cases (either $O_{p}(1)$ or vanishing), and its behavior will not change when we use the full set of averages $\widehat{\mathbf{F}}=[\overline{\mathbf{y}},\overline{\mathbf{X}}]$. Also, $\left\|\sqrt{NT}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|=O_{p}(1)$ under general factors (see Westerlund,, 2018), and thus the $\sqrt{NT}$-consistency will hold under our assumptions, as well. Similarly,
@@
$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})\right\|$ $\leq(NT)^{-1/2}\left\|\sqrt{NT}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$
$\times\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})\right\|$
$=\begin{cases}O_{p}((NT)^{-1/2}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.87)
@@
because asymptotically $\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$ behaves identically to $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$, and thus the dominating component will have the same rate. Eventually, by following the same analysis of $\mathbf{I},\mathbf{II}$ and $\mathbf{III}$ in (5.3.1), we have
$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})\right\|=\begin{cases}O_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ O_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M.\end{cases}$ (A.88)
In fact, we can demonstrate that in analogy to the analysis of (5.3.1), we can decompose and simplify the term:
$\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})-\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
$+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}(\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})+o_{p}(T^{-\tau})$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}(\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i})+o_{p}(T^{-\tau})$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\gamma}_{i}^{\prime}\mathbf{F}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}\boldsymbol{\gamma}_{i}+o_{p}(T^{-\tau})$
$=c_{NT}+o_{p}(T^{-\tau})$ (A.89)
for a strictly positive $c_{NT}$ with $c_{NT}\to c>0$ under $M\subset M_{0}$.
To complete, we will assume that $\tau\in(0,0.5)$, in order to isolate $o_{p}(T^{-\tau})$ terms. Then we have
$\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}=c_{NT}+o_{p}(T^{-\tau}),$ (A.90)
and, because we can show that $b_{NT}:=\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}\widehat{\boldsymbol{\nu}}_{i}$ converges to a positive scalar quantity, we can analogously show that
$(\mathrm{IC}^{MW}(M)$ $-\mathrm{IC}^{MW}(M_{0}))/\ln(T)=\ln(1+T^{\tau}(c_{NT}\times b_{NT}^{-1}+o_{p}(T^{-\tau}))/\ln(T)+o(1)$
$=\ln(1+T^{\tau}c_{NT}\times b_{NT}^{-1})/\ln(T)+o_{p}(1)$
$=\ln(T^{\tau}c\times b_{NT}^{-1})/\ln(T)+\left(\ln(1+T^{\tau}c_{NT}\times b_{NT}^{-1})-\ln(T^{\tau}c_{NT}\times b_{NT}^{-1}))\right)/\ln(T)$
$+o_{p}(1)$
$=\ln(T^{\tau}c_{NT}\times b_{NT}^{-1})/\ln(T)+\frac{1}{\ln(T)}\ln\left(1+\frac{b_{NT}}{cT^{\tau}}\right)+o_{p}(1)$
$=\tau+\ln(c_{NT}\times b_{NT}^{-1})/\ln(T)+\frac{1}{\ln(T)}\left(\frac{b_{NT}}{c_{NT}T^{\tau}}+O(T^{-2\tau})\right)+o_{p}(1)$
$\to_{p}\tau>0$ (A.91)
as $(N,T)\to\infty$. Therefore,
$\mathbb{P}(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})<0)=\mathbb{P}\left(\frac{\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})}{\ln(T)}<0\right)\to 0$ (A.92)
as $(N,T)\to\infty$. Under $M_{0}\subset M$, the proof goes exactly the same as in Margaritella and Westerlund, (2023),and so given the rates declared above and the fact that $p_{N,T}^{-1}O_{p}(C_{N,T}^{-2})=o(1)$, we obtain
$p_{N,T}^{-1}(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0}))=(g-m)+o_{p}(1)$ (A.93)
and so
$\mathbb{P}\left(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})<0\right)=\mathbb{P}\left(p_{N,T}^{-1}(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0}))<0)\right)\to 0,$ (A.94)
and the overspecification risk is eliminated.
5.3.3.2  Heterogeneous $\boldsymbol{\beta}$